OpenStudy (anonymous):

if sin A=3/5 and cos B=-12/13, A and B obtuse, find: a) sin(A-b) b)cos(A+B)

OpenStudy (campbell_st):

well you 1st need to find the missing sides in each triangle.. both 2nd quadrant angles  the use sin(A - B) = sinAcos(B) - cos(A)sin(B) and substitute the exact values from the triangles. and for (b) cos(A + B) = cos(A)cos(B) - sin(A)sin(B) again substitute and simplify hope it helps

OpenStudy (anonymous):

I'll try it out thank you

OpenStudy (anonymous):

I got A.) -16/65, B.) 33/65

OpenStudy (anonymous):

A.) sin(3/5)cos(-12/13)-cos(-4/5)sin(5/13) B.) cos(-4/5)cos(-12/13)-sin(-3/5)sin(5/13)